### v1

# AI Feedback Enhances Community-Based Content Moderation through Engagement with Counterarguments

Abstract
Today, social media platforms are significant sources of news and political communication, but their role in spreading misinformation has raised significant concerns. In response, these platforms have implemented various content moderation strategies. One such method, Community Notes on X, relies on crowdsourced fact-checking and has gained traction, though it faces challenges such as partisan bias and delays in verification. This study explores an AI-assisted hybrid moderation framework in which participants receive AI-generated feedback —supportive, neutral, or argumentative —on their notes and are asked to revise them accordingly. The results show that incorporating feedback improves the quality of notes, with the most substantial gains resulting from argumentative feedback. This underscores the value of diverse perspectives and direct engagement in human-AI collective intelligence. The research contributes to ongoing discussions about AI’s role in political content moderation, highlighting the potential of generative AI and the importance of informed design.
Keywords— Content Moderation; Human-AI Interaction; Large Language Models; Polarisation; Hybrid Intelligence
Introduction
The emergence of social media platforms has transformed how users consume and share information. Studies show that these platforms have become primary news sources for many individuals [1, 2] and are widely used by politicians as campaign tools [3]. This shift presents challenges for platform owners in monitoring and verifying the information being shared. The widespread dissemination of misleading content and misinformation has become a primary concern for social media platforms. This issue is especially pronounced during significant socio-political events, such as the 2016 U.S. presidential election [4] and the Brexit referendum [5]. The spread of false information during times of crisis can seriously mislead the public. For example, the rapid circulation of misinformation about COVID-19 led the WHO’s Director-General to describe the situation as an ”infodemic” [6]. Another striking case is the January 6th, 2021, attack on the U.S. Capitol, where social media posts helped spread the misinformation that incited violence [7]. In response to these concerns, social media platforms have implemented various content moderation strategies, including both community-based and automated approaches.
On January 7, 2025, Mark Zuckerberg announced the end of Meta Platforms’ third-party fact-checking program, replacing it with Community Notes [8]. Launched initially on X (formerly Twitter) in January 2021, Community Notes represented the first large-scale, community-based initiative to moderate misleading content [9]. This shift in moderation strategy reflects the limitations of traditional fact-checking, which relies on human experts to review and classify large volumes of content. This approach is both expensive and challenging to scale [10]. Moreover, expert-driven moderation has faced criticism for potential political and cultural bias. A 2020 report found that 90% of Republicans in the United States believed social media platforms censor content based on political views [11].
Community Notes rely on the wisdom of the crowd to identify misleading content and leverage the diversity of users on the platform to produce moderated outcomes. The process involves displaying specific posts to users in the Community Notes program and asking them to write a note that provides context for those posts. Users are also asked to indicate whether they believe the post is misleading and to provide reasoning for their judgment. Subsequently, other users rate the helpfulness of these notes. Initially, notes with high helpfulness ratings appeared beneath the posts in users’ timelines. However, early analyses revealed strong partisan behaviour. Yasseri and Menczer (2023) found that the network connecting note writers and raters mirrors the highly polarised friendship network on X [12]. Similarly, Allen et al. (2022) showed that users are more likely to write notes on posts by counter-partisans, label them as misleading, and rate notes from counter-partisans as unhelpful, while rating notes from co-partisans as helpful [13].
To address partisan behaviour among Community Notes users, adjustments were made to the rating algorithm. Notes that receive helpful ratings from users with diverse perspectives, who typically rate content differently, are now prioritised over those endorsed only by ideologically similar users [9]. Although this change was intended to reduce polarised interactions, it has increased the time required for notes to be selected and attached to posts. A recent report found that 74% of accurate Community Notes addressing misinformation about the 2024 U.S. presidential election never reached users, allowing misleading content to go unchallenged. As a result, misinformation posts about the election spread 13 times faster than those accompanied by fact-checked notes [14]. These delays undermine the moderation process, as research shows that lag in content moderation significantly reduces its effectiveness [15].
One approach to addressing the challenges of scale and speed in moderation is the use of automated algorithms. Early efforts in this area have focused on applying machine learning to assess the truthfulness of statements [16]. These systems typically rely on supervised learning and require pre-labelled datasets, such as LIAR [17], which compiles fact-checked claims from dedicated websites, and FEVER [18], which includes claims derived from Wikipedia. However, building high-performing models remains difficult due to the need for large, diverse labelled datasets—a process that demands significant labour from experts or crowdworkers [19]. The effectiveness of these automated systems is also limited by the quality and biases of their training data, which can perpetuate inaccuracies in classification [20].
However, recent advances in generative AI offer promising solutions to the limitations of existing automated content moderation strategies [21, 22]. Research shows that Large Language Models (LLMs) can effectively persuade individuals on political issues [23] and generate fact-checking messages that are more persuasive than those produced by human experts [24]. In addition, several studies demonstrate that LLMs can accurately assess the credibility of news sources [25] and verify factual claims [26, 27, 28], with their accuracy improving when fine-tuned on augmented data [29]. In a recent example, De et al. (2025) fine-tuned an LLM on the Community Notes dataset to generate what they call “supernotes,” which were rated significantly more helpful than notes written by Community Notes users for the same posts [30]. These findings suggest that LLMs are not only capable of understanding complex content but may also outperform human-written moderation in certain cases. However, studies have also found that purely automated approaches can be less effective in the presence of social influence [31]. Moreover, people tend to place considerable trust in LLMs, even when the models make mistakes, raising concerns about overreliance on AI-generated content [32].
An increasingly supported view in the literature is that the most effective content moderation strategies combine human oversight with AI assistance [33]. This hybrid approach leverages the complementary strengths of humans and machines, enabling moderation that is both more nuanced and scalable [34]. Schmitt et al. (2024) suggest that models integrating human judgment with AI can outperform either component alone [35]. However, even within hybrid systems, users often overestimate the accuracy of LLMs, placing trust in AI-generated assessments even when they are incorrect—a phenomenon known as automation bias or machine heuristics [36, 37]. Moreover, some Human-AI frameworks risk reinforcing users’ existing beliefs, potentially deepening echo chambers and contributing to political polarisation [38]. To mitigate these risks, such systems must be carefully designed to avoid unintentionally amplifying bias and division on social media platforms.
Despite these risks, political diversity itself can be a valuable asset in improving community-based moderation efforts. One of the key strengths of Community Notes is its ability to harness ideological diversity—even polarisation—to produce more balanced and neutral notes [39]. A similar dynamic has been observed on Wikipedia, where research attributes the platform’s success to the variety of viewpoints among editors, along with its structured policies and guidelines [40, 41, 42, 43]. Shi et al. (2019), for instance, found that higher levels of polarisation among editors correlate with higher-quality articles, especially on politically sensitive topics [42]. Moreover, studies suggest that exposure to cross-partisan messages can reduce both affective and issue-based polarisation [44, 45]. These findings indicate that, when properly structured, political diversity not only does not hinder collaboration but also can even enhance it. In a recent study, researchers paired two participants to collaboratively write Community Notes. They found that duos produced more helpful notes than individuals, and that politically diverse groups were especially effective when evaluating posts with Republican-leaning content [46]. These results further support the idea that political diversity can strengthen the note-writing process in Community Notes, provided it is applied in a way that is both effective and efficient.
Building on these insights, this study proposes a human-AI collaboration framework that uses AI-generated feedback to simulate political diversity in content moderation. The aim is to combine the efficiency of AI with the interpretive strengths of human contributors, while ensuring that humans retain final decision-making authority. We investigate whether AI-generated feedback can emulate cross-partisan collaboration and enhance the quality of contributions on the Community Notes platform. Our focus is on how users revise their notes in response to different types of AI-generated feedback. Specifically, this study addresses the following research questions:
RQ1: How does the perceived quality of notes vary by the type of feedback received (argumentative, supportive, or neutral)?
RQ2: How does the impact of feedback on perceived note quality differ when participants evaluate co-partisan versus cross-partisan posts?
RQ3: How does the effect of feedback on perceived quality change based on whether participants believe it was generated by a human expert or an AI model?
To investigate these research questions, we conducted an experiment modelled after the Community Notes process. Participants () were shown a post and asked to write a note providing additional context. Each note was then submitted to a GPT-4 language model, which was prompted to generate one of three types of feedback: argumentative, supportive, or neutral.
Argumentative feedback offers counterpoints to the original note, supportive feedback reinforces the note’s arguments, and neutral feedback simply restates its content. While all feedback was generated by the AI model, we randomly told some participants that the feedback was written by a human expert. This manipulation allows us to examine the roles of human favouritism [47] and automation bias [36] in shaping participants’ responses.
Participants were then asked to revise their notes in light of the feedback. In a second experiment (), we crowdsourced evaluations of note quality before and after revision. Self-identified Republicans and Democrats rated the notes on a 0–10 scale based on their helpfulness. Figure 1 provides a schematic overview of the experimental workflow.
Results
Participants engaged with AI-generated feedback to varying degrees. Figure 2 presents two contrasting cases in which participants received argumentative feedback from the AI. In Figure 2(a), the participant maintained their original stance but incorporated phrasing from the feedback into the revised note. In contrast, in Figure 2(b), the participant disregarded the feedback entirely, making no meaningful changes.
Comparing how these revised notes were evaluated by Democrat and Republican raters reveals a clear pattern. The note in Figure 2(a), which engaged with the feedback, received higher ratings. Conversely, the note in Figure 2(b), which ignored the feedback, was rated poorly.
These contrasting outcomes highlight a key insight: the extent to which participants engage with AI-generated feedback significantly influences the effectiveness of their revisions.
To measure engagement, we introduce the Feedback Acceptance rate (), a metric that quantifies the extent to which participants incorporated the provided feedback into their final submissions. Specifically, is defined as the semantic similarity score between the feedback and the revised note (see Methods for details). To evaluate whether engagement with feedback leads to higher-quality revisions, we use as a predictor in ordinal logistic regression models. The dependent variables capture whether the revised note has declined, was unchanged, or improved according to the Democrat raters () and Republican raters (). The results (Table 1) show that is a strong and consistent predictor of note improvement as perceived by both partisan groups. A comparison of odds ratios further indicates that has a slightly stronger effect on improvement ratings from Democrats than from Republicans.
| Ordinal Logistical Regression | Dependent variable: | Dependent variable: | ||
| odds ratio | odds ratio | |||
| 3.581(0.604) | <0.001 | 2.498(599) | <0.001 | |
| Observations | 893 | 893 | ||
| Pseudo | 0.019 | 0.009 | ||
| Deviance | 1872.86 | 1924.63 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -936.43 | -962.32 | ||
| Akaike Inf. Crit. | 1878.86 | 1930.63 | ||
To better understand the factors that influence this engagement metric, we analyse the relationship between and three key variables: the type of post (supporting either Democratic or Republican ideology), the participant’s self-identified partisanship, and the type of feedback received. The results from the Ordinary Least Squares (OLS) regression are shown in Table 2.
The analysis shows that users who received argumentative feedback were less likely to incorporate it into their final note than those who received supportive feedback. Additionally, Democratic participants were more likely than Republicans to use the feedback in their final notes. However, the type of post did not have a significant effect on this behaviour.
However, analysing the type of post in isolation may not provide the complete picture. We expect individuals who share the political ideology of the post’s author to respond differently when moderating content compared to those with opposing views. To investigate this, we introduce a binary alignment variable that captures the relationship between a participant’s self-identified partisanship and the post’s stance: co-partisan if the participant and the post share the same ideology, and cross-partisan if they do not. A post is classified as co-partisan when it supports Democrats (Republicans) and the participant identifies as a Democrat (Republican); otherwise, it is classified as cross-partisan.
| Ordinary Least Squares (OLS) Regression | Dependent variable: | |
| Coefficient | ||
| Feedback (Argue) | -0.094(0.008) | <0.001 |
| Feedback (Support) | -0.052 (0.008) | <0.001 |
| Post (D) | 0.001 (0.007) | 0.880 |
| Participant (D) | 0.021 (0.007) | 0.001 |
| Constant | 0.850 (0.007) | <0.001 |
| Observations | 893 | |
| 0.143 | ||
| Adjusted | 0.139 | |
| Residual Std. Error | 0.097 (df = 888) | |
| F Statistics | 36.899 (df = 4; 888) | |
To examine how political alignment affects engagement with feedback, we conducted an OLS regression with as the dependent variable, including an interaction term between co-partisanship and feedback type. The results of this analysis are shown in Table 3. The findings indicate that when participants share the same political affiliation as the post, they are less likely to incorporate argumentative feedback. A similar pattern emerges for supportive feedback: participants are also less likely to adopt supportive feedback language when the post aligns with their political views.
| Ordinary Least Squares (OLS) Regression | Dependent variable | |
| Coefficient | ||
| Co-partisan × Neutral | -0.019 (0.011) | 0.090 |
| Cross-partisan × Argue | -0.102 (0.012) | <0.001 |
| Co-partisan × Argue | -0.104 (0.011) | <0.001 |
| Cross-partisan × Support | -0.059 (0.011) | <0.001 |
| Co-partisan × Support | -0.064 (0.011) | <0.001 |
| Constant | 0.871 (0.008) | <0.001 |
| Observations | 893 | |
| 0.135 | ||
| Adjusted | 0.130 | |
| Residual Std. Error | 0.098 (df = 887) | |
| F Statistics | 27.711 (df = 5; 887) | |
The second treatment condition manipulated the labelled source of the feedback. Specifically, we investigate whether participants respond differently depending on whether they believe the feedback was generated by a human expert or an AI agent. To explore this, we replicate the previous regression analysis on two separate subsets of the data: one in which feedback was attributed to an AI agent, and another in which it was attributed to a human expert. Figure 3 presents the marginal predicted values of across the different treatment conditions for both feedback source labels. While the overall patterns are consistent across conditions, we observe a slight increase in when the feedback is labelled as AI-generated, specifically for cross-partisan posts. This pattern holds for both argumentative and supportive feedback. These results suggest a marginal preference for AI-labelled feedback when participants engage with content that opposes their political views. However, the differences are not statistically significant (e.g., predicted for AI in cross-partisan × argumentative = 0.777 vs. 0.762 for human expert, ; predicted for AI in cross-partisan × supportive = 0.822 vs. 0.803 for human expert, ), indicating that the labelled source of feedback does not meaningfully influence user behaviour in this context.
To construct the full model and evaluate the effect of feedback type conditional on participant engagement, we estimated an ordinal logistic regression including both the main effects of feedback type and its interaction with . The dependent variables are and , which reflect the improvement of the notes according to Democrat and Republican raters, respectively. This model allows us to assess whether argumentative or supportive feedback is more likely to lead to improvements in perceived helpfulness when participants actively engage with the feedback, and whether these effects differ across political subgroups.
The results of this analysis are presented in Table 4. The findings indicate that once participants incorporate feedback into their notes, both argumentative and supportive feedback increase the likelihood of producing a more helpful note, as rated by both Democratic and Republican evaluators. However, argumentative feedback shows a larger effect size than supportive feedback, suggesting that when participants engage with feedback that challenges their views and includes counterarguments, the resulting notes are more likely to be of higher quality.
| Ordinal Logistical Regression | Dependent variable | Dependent variable | ||
| odds ratio | odds ratio | |||
| Feedback (Argue) | -1.210 (1.362) | 0.374 | -1.944 (1.346) | 0.148 |
| Feedback (Support) | -0.583 (1.228) | 0.635 | -0.765 (1.238) | 0.536 |
| Feedback (Neutral) × | 2.643 (0.987) | 0.007 | 1.534 (1.001) | 0.125 |
| Feedback (Argue) × | 3.713 (1.370) | 0.006 | 3.817 (1.331) | 0.004 |
| Feedback (Support) × | 3.662 (1.086) | <0.001 | 2.237 (1.079) | 0.038 |
| Observations | 893 | 893 | ||
| Pseudo | 0.027 | 0.011 | ||
| Deviance | 1857.37 | 1921.71 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -928.68 | -960.86 | ||
| Akaike Inf. Crit. | 1871.37 | 1935.71 | ||
Discussion
As Community Notes gains traction across major social media platforms as a leading approach to crowd-based content moderation [8], concerns have been raised about its efficiency and long-term sustainability [13, 48, 15]. One emerging response to these challenges is the integration of generative AI—particularly large language models (LLMs)—into the Community Notes ecosystem. In this study, we introduce a framework that leverages GPT-4 to provide users with different types of feedback on their notes. This framework is designed both to evaluate the effectiveness of AI-assisted feedback and to deepen our understanding of the dynamics of human–AI collaboration in this context. We found that within this framework, the effectiveness of feedback in improving note quality depends not only on its type but also on the extent to which participants engage with it.
Studies have shown that two key factors in collective intelligence are the effort a group invests and how effectively members leverage each other’s knowledge and skills [49]. Our findings align with this: participants who engaged with AI-generated feedback more and incorporated it into their revisions produced better perceived notes. Specifically, the Feedback Acceptance rate ()—which reflects how closely participants’ final notes match the feedback—emerged as a significant predictor of note improvement for both Democratic and Republican raters ( and ). A high indicates strong engagement with the feedback, while a low often corresponds to users disregarding the feedback entirely, sometimes resulting in notes of lower quality than the original (See Figure SI10). This behaviour mirrors patterns observed on social media, where users often react defensively to content that challenges their political beliefs [50]. Our findings suggest that when participants move beyond this reflexive partisan response and actively engage with feedback, they produce higher-quality notes.
Among participants who actively engaged with the feedback, our findings show that argumentative feedback was more likely to lead to improvements in note quality. This may be because engaging with counterarguments encourages users to reconsider and refine their positions, resulting in notes that are more balanced, better justified, and ultimately more helpful. Prior research suggests that in problem-solving contexts, when AI systems are designed to complement human cognitive biases, they can prompt users to discover novel solutions, thereby improving outcomes [51]. Relatedly, Juncosa et al. [46] demonstrate that diversity among groups of note authors can enhance note quality under certain conditions. Our current findings suggest that argumentative AI feedback may function as a form of political diversity, exposing users to alternative viewpoints and, in doing so, fostering the production of higher-quality content.
In our second research question, we examine how the political alignment between participants and posts affects the perceived helpfulness of the notes. The results indicate that when participants and posts share the same political ideology (i.e., are co-partisan), participants are less likely to incorporate feedback into their notes, reducing the likelihood of producing a higher-quality note. This pattern holds for both argumentative and supportive feedback. One possible explanation for this behaviour is that when participants encounter a post aligned with their own beliefs, they may become less open to revision, regardless of the content or tone of the feedback [52, 53].
Lastly, in our third research question, we examined how the labelled source of feedback, whether an AI agent or a human expert, influences participant behaviour and, consequently, the quality of their notes. We find a slight preference for AI-labelled feedback when the post comes from the opposing political party (i.e., in cross-partisan contexts). In these cases, participants are more likely to engage with feedback they believe was generated by an AI agent. This finding is consistent with prior research on automated content moderation, which suggests that users tend to be more receptive to algorithmic decisions. This tendency, known as automation bias, reflects the inclination to overestimate the neutrality and competence of automated systems, even when they make errors [36, 32].
This study offers insights into the evolving role of AI in crowd-based content moderation. We demonstrate that integrating a large language model (LLM) into the Community Notes system can enhance its efficiency, provided the design preserves human agency and the platform’s core participatory nature. In our framework, the AI functions solely as a reviewer offering suggestions, while users retain full control over the content of their notes, helping to prevent the imposition of any algorithmic bias.
A key contribution of our approach is the introduction of political diversity at the note-writing stage, rather than relying solely on diverse raters at the evaluation stage. When participants engage with argumentative feedback, they are more likely to produce higher-quality notes, as rated by both Democratic and Republican evaluators. This suggests that such feedback not only promotes quicker production of helpful notes but also encourages users to consider counterarguments and revise their contributions in a more balanced and thoughtful way, ultimately discouraging the reinforcement of partisan narratives.
While our findings highlight the potential of integrating argumentative AI feedback into Community Notes, there are important limitations to consider. First, we do not assess the objective truthfulness of the notes. Instead, note quality in this study is evaluated through crowdsourced ratings, with groups of Democrats and Republicans judging notes based on their perceived helpfulness. This method does not provide a direct measure of factual accuracy—for instance, a professional fact-checker might reach different conclusions. However, our approach reflects the current practice in Community Notes, where community members determine which notes are approved and displayed beneath posts. While some researchers question the reliability of crowd-based methods for identifying misinformation, it is important to note that the goal of Community Notes is not necessarily to debunk falsehoods, but to add context to potentially misleading posts. These posts may not contain explicit inaccuracies, but can still lead readers to incorrect conclusions due to missing or unbalanced information. In such cases, a politically diverse and balanced group of laypeople may be well positioned to assess whether the added context is informative and helpful.
Second, findings from an experimental setting may not fully generalise to real-world contexts. For example, participants in Community Notes are volunteers who may be motivated by political or ideological beliefs, whereas participants in our study were recruited and compensated for their involvement. Third, the “breaking news problem” remains a persistent challenge. For AI to provide meaningful feedback, it must have a sufficient understanding of the topic. When this is lacking, the AI may still offer feedback that encourages critical thinking, effectively acting as a nudging algorithm [54]. That said, recent advancements in large language models have enabled web-search capabilities, which may allow them to engage more effectively with current events in future applications.
Overall, this study provides a deeper understanding of human–AI interaction within a community-based content moderation system. Our findings suggest that AI can meaningfully enhance this process, not by replacing human judgment, but by offering targeted feedback. In particular, AI-generated argumentative feedback can introduce political diversity and improve note quality by prompting authors to consider opposing viewpoints. However, the presence of AI alone is insufficient to ensure effective collaboration. Its impact depends on active user engagement, which is shaped by factors such as the participant’s partisanship and the political alignment of the post.
Future work could extend these insights by examining longer-term interactions between users and AI to understand how engagement patterns evolve over time. Allowing users to interact with the AI across multiple sessions would provide a deeper understanding of sustained behaviour in this context. Additionally, while this study focused on political content, it would be valuable to explore how human–AI collaboration functions in apolitical or less polarising domains to assess whether similar engagement patterns emerge.
Methods
Data Collection
We collected data for this study through a series of experiments involving participants recruited via the online research platform Prolific. Eligibility was limited to residents of the United States. Participants received compensation for their time. The experiments were designed and implemented using the oTree framework (version 5.10.4), and GPT-4 was used as the AI agent in the note-writing task. In this section, we provide a detailed overview of the experimental procedures, including the creation of notes and their subsequent evaluation through crowd-based ratings.
Note Writing
In this experiment, participants first indicated their political affiliation using a nine-point slider ranging from ”Strongly Democrat” to ”Strongly Republican.” Those who identified as neutral were excluded from the study. Participants were then shown a political social media post and asked to write a contextual note and classify the post as either misleading or not. The instructions mirrored those provided to Community Notes users. After submitting their initial notes, participants received feedback labelled as coming from either an AI agent or a human expert, though in all cases, the feedback was generated by GPT-4. Participants were debriefed about the true source of the feedback at the end of the experiment.
Participants were randomly assigned to one of three feedback conditions. In the Argue group, the AI provided feedback that challenged the points made in the participant’s note while maintaining an opposing stance. In the Neutral group, the AI rephrased the note without introducing contradictions or additional evidence. In the Support group, the AI reinforced and expanded on the participants’ original arguments. Examples of AI-generated feedback for each condition are shown in Figure 4, and the full prompts used to generate this feedback are included in the Supporting Information.
After receiving feedback, participants were asked to revise their notes and were given the option to update their original classification of the post as misleading or not. Finally, they completed a short survey measuring their comfort level with emerging technologies. Additional details about the experiment are provided in the Supporting Information.
We used 40 social media posts with clear partisan messaging, all previously analysed by Community Notes users. After data cleaning, the final dataset included 893 participants: 455 self-identified Democrats and 438 self-identified Republicans. The study received ethical approval from the University College Dublin Office of Research Ethics (HS-LR-24-230-Mohammadi-Yasseri).
Note Evaluation
In the next phase of data collection, we recruited a new group of participants to evaluate the notes. After providing informed consent, participants first reported their political partisanship using a nine-point scale. They were then shown two different social media posts, each accompanied by 10 notes, and asked to evaluate the notes along two dimensions: helpfulness and favouritism. Helpfulness was rated on a scale from 0 to 10, reflecting the perceived utility of the note. Favouritism was assessed using a nine-point scale, indicating the extent to which a note appeared biased toward either political party, ranging from ”Favouring Democrats” to ”Favouring Republicans.” As shown in Figure SI11, each note received evaluations from at least three Democratic and three Republican participants. In total, we recruited 1354 participants for this evaluation phase, with each participant evaluating 20 notes across two posts (10 notes per post). This part of the study too received ethical approval from the University College Dublin Office of Research Ethics (HS-LR-24-188-Mohammadi-Yasseri).
Measurements
Helpfulness
The data collected from the note evaluation experiment provides the helpfulness scores assigned to each note by participants identified as Democrats and Republicans. As described earlier, each evaluator rates two posts and 10 notes per post. To standardise the scores and account for the rating habit of each user, we calculated z-scores of the notes associated with each post by a rater. The normalised score , representing helpfulness given by user to note for post , is calculated as
where is the mean of the scores assigned by user to post , and is the standard deviation of these scores.
To ensure reliability, we excluded the highest and lowest ratings for each note, provided the note had been rated by at least five Democrats and five Republicans (evaluated separately). We then calculate the final helpfulness score for each note by averaging the remaining normalised ratings.
Improvement is calculated as the relative percentage change in helpfulness scores from the initial to the final note. The improvement metric , where indicates Democrat or Republican raters, is defined as:
Notes were then categorised based on the percentage change in improvement. Notes with an improvement greater than 10% are classified as improved (). Those with a change between and are classified as unchanged (). Notes with a decline greater than 10% are labelled as declined ().
Feedback Acceptance rate ()
To quantify the extent to which participants incorporated the feedback into their final notes, we calculated the semantic similarity between each piece of feedback and the corresponding revised note. This was measured using cosine similarity, implemented via the Scikit-learn package (version 0.0.post9). The resulting metric, referred to as the Feedback Acceptance rate (), captures the degree to which the feedback has been incorporated in the revised note. Figure SI8 displays the distribution of scores across the different feedback conditions.
The same semantic similarity method was also used to compute the similarity between the initial and revised versions of each note. This measure captures the extent to which a note changed during the revision process. Lower similarity scores indicate greater changes. Figure SI9 shows the distribution of similarity scores between the initial and final notes across the different feedback conditions.
Statistical Analysis
We conducted ordinal logistic regressions to examine changes in helpfulness scores, which could take values of 1 (”improved”), 0 (”unchanged”), or -1 (”declined”). The regressions were performed using the polr() function from the MASS package (version 7.3.57) in R. We report odds ratios, standard errors, and two-tailed p-values for each predictor. The p-values were calculated using the Wald test. To investigate the variables affecting the Feedback Acceptance rate (), we used ordinary least squares regressions (OLS). We used the lm() function from the stats package (version 4.2.1) in R to estimate the models. We report coefficient estimates, standard errors, and two-tailed p-values for each predictor. The p-values were calculated using the Wald test.
Data Availability
Data used in the study are available on Harvard Dataverse (doi:10.7910/DVN/SWTACT).
Acknowledgments
The authors wish to thank Gabriela Juncosa and Nico Mutzner for their valuable comments. This publication has emanated from research supported in part by a grant from Taighde Éireann – Research Ireland under Grant numbers 18/CRT/6049 and IRCLA/2022/3217. TY acknowledges support from Workday, Inc. For the purpose of Open Access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.
Author Contribution
SM and TY designed the experiments and analysed the data. SM conducted the experiments. SM and TY drafted the paper. TY secured the funding and supervised the project. All authors contributed to writing the manuscript and gave final approval for publication.
Competing interests
The authors declare no competing interests.
Correspondence
Correspondence and requests for materials should be addressed to Taha Yasseri: taha.yasseri@tcd.ie.
Supplementary Information for
AI Feedback Enhances Community-Based Content Moderation through Engagement with Counterarguments
Saeedeh Mohammadi and Taha Yasseri
Note Writing Experiment
In this experiment, participants first viewed an information sheet detailing the risks and benefits of participation, the study’s purpose, and ethics approval information. After providing consent and entering their Prolific ID, participants were asked to indicate their political partisanship on a nine-point scale ranging from strongly Democrat to strongly Republican. A screenshot of this page is shown in Figure SI1. Participants who selected the neutral midpoint option were excluded from the experiment.
Afterwards, participants were presented with a description of the study. At this stage, they were informed that the feedback they would receive could come either from a human expert or an artificial intelligence machine. Figure SI2 shows a screenshot of this page.
In the first task of the experiment, participants were instructed to read a social media post and write a note aimed at providing additional context and assessing whether the post was misleading or not. The note was required to be under 280 characters. Participants were given 10 minutes to complete this task. Figure SI3 shows an example of the task page.
After completing the first task, participants were shown a “please wait” page, which informed them about the source of the upcoming feedback. They remained on this page for 2.5 minutes to create the impression that the feedback might be provided by a human. Figure SI4 displays two screenshots of this page, illustrating both feedback source conditions.
Afterwards, participants were given seven minutes to review the feedback and revise their notes accordingly. During this stage, they were also allowed to change the classification of the post (misleading or not misleading). Additionally, participants were reminded of the feedback source. An example of this page is shown in Figure SI5.
After the second task, participants completed a post-experiment survey where they rated their comfort level with various technologies. Figure SI6 shows the survey page displaying the full list of questions. Participants were given four minutes to complete the survey.
Subsequently, participants were asked to identify the source of the feedback as an attention check. Those who selected an incorrect option or chose “I do not remember” were excluded from the dataset. Finally, on the debrief page, participants were informed that all feedback was generated by an AI agent and that no human experts were involved.
Note Evaluation Experiment
In the note evaluation experiment, participants were recruited to assess each note produced in the note writing experiment. Upon entering the study, participants first viewed an information sheet outlining the risks and benefits of participation, the purpose of the study, and ethics approval details. They then completed a consent form and provided their Prolific ID.
Afterwards, participants were asked to indicate their partisanship on a nine-point scale, consistent with the note-writing experiment. Those who selected the “neutral” option were excluded from the study.
In the main task of the experiment, participants were shown a social media post accompanied by 10 notes written about that post. They were asked to rate each note on two dimensions: overall helpfulness and favouritism. Overall helpfulness was defined as the quality and usefulness of the note in providing appropriate context for the post, rated on a scale from 0 to 10. Favouritism was described as the degree to which the note displayed bias favouring one political party beyond objectivity, measured on a nine-point scale ranging from strongly favouring Democrats to strongly favouring Republicans. Participants had 7 minutes to evaluate all 10 notes. Figure SI7 shows a screenshot of the task description. Each participant evaluated two different social media posts, each with its own set of 10 notes.
Protocol for selecting the Social Media Posts
We selected 40 different posts from X.com for the note-writing experiment. These posts were drawn from the Community Notes dataset as of the end of August 2024. The selection followed the criteria below:
The post contains at least one of the following keywords: Democrat, Republican, Biden, or Trump.
It has at least two notes tagged as ”misinformed” in the Community Notes dataset.
The post is related to the United States.
The text is clear and easy to understand.
The post does not include a video or a link to an article required to understand its content.
Afterwards, the posts were classified as either pro-Democrat or pro-Republican based on their content. In total, we collected 20 pro-Democrat posts and 20 pro-Republican posts.
Feedback Prompts
We used the OpenAI API (version 1.25.0) to prompt the GPT-4 model for generating feedback in the note-writing experiment. The ”chat.completion” function was employed, with the system message set as: ”You are a helpful assistant who writes a note based on the instruction. The note should be limited to 280 characters. Do not use any hashtags or links in your response.” The temperature parameter was set to 1. The user message varied depending on the type of feedback. The specific prompts used are as follows:
Argue: Read the following tweet:tweet. Read the comment made in response to the tweet:comment. Argue with the comment while maintaining a consistent opposing stance.
Support: Read the following tweet:tweet. Read the comment made in response to the tweet: comment. Support the comment’s point. Help justify this comment, add more arguments if you can.
Neutral: Read the following tweet:tweet. Read the comment made in response to the tweet: comment. Rephrase the comment. Do not add any evidence or contradict the comment in any way.
The ”tweet” refers to the content of the post, while the ”comment” is created by appending the sentence “The tweet is misleading” or “The tweet is not misleading” to the note written by the participant in the first task. This sentence is selected based on the participant’s classification of the post. The model is explicitly prompted to avoid including links, reducing the risk of hallucinated content generated by the AI.
Additional Figures
Additional Tables
| Ordinal Logistical Regression | Dependent variable | Dependent variable | ||
| odds ratio | odds ratio | |||
| Post (D) × Participant (R) × Feedback (Neutral) | 0.449 (0.298) | 0.132 | 0.294 (0.307) | 0.338 |
| Post (R) × Participant (D) × Feedback (Neutral) | 0.424 (0.299) | 0.156 | -0.202 (0.298) | 4.9801 |
| Post (D) × Participant (D) × Feedback (Neutral) | 0.102 (0.307) | 0.739 | -0.318 (0.301) | 0.291 |
| Post (R) × Participant (R) × Feedback (Argue) | -0.676 (0.304) | 0.026 | -0.292 (0.299) | 0.328 |
| Post (D) × Participant (R) × Feedback (Argue) | -0.570 (0.313) | 0.068 | -0.317 (0.301) | 0.293 |
| Post (R) × Participant (D) × Feedback (Argue) | -0.189 (0.306) | 0.537 | -0.403 (0.309) | 0.192 |
| Post (D) × Participant (D) × Feedback (Argue) | -0.162 (0.295) | 0.583 | -0.569 (0.294) | 0.0533 |
| Post (R) × Participant (R) × Feedback (Support) | 0.249 (0.302) | 0.409 | 0.240 (0.308) | 0.436 |
| Post (D) × Participant (R) × Feedback (Support) | 0.144 (0.301) | 0.631 | -0.752 (0.297) | 0.011 |
| Post (R) × Participant (D) × Feedback (Support) | 0.644 (0.302) | 0.033 | -0.488 (0.298) | 0.133 |
| Post (D) × Participant (D) × Feedback (Support) | 0.277 (0.301) | 0.358 | -0.329 (0.297) | 0.268 |
| Observations | 893 | 893 | ||
| Pseudo | 0.019 | 0.011 | ||
| Deviance | 1872.58 | 1920.29 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -936.29 | -960.15 | ||
| Akaike Inf. Crit. | 1898.58 | 1946.29 | ||
| Ordinal Logistical Regression | Dependent variable | Dependent variable | ||
| odds ratio | odds ratio | |||
| Post (D) | -0.025 (0.125) | 0.840 | 0.145 (0.124) | 0.242 |
| Participant (D) | 0.251 (0.125) | 0.045 | -0.225 (0.124) | 0.069 |
| Feedback (Argue) | -0.644 (0.155) | <0.001 | -0.332 (0.152) | 0.029 |
| Feedback (Support) | 0.077 (0.152) | 0.612 | -0.268 (0.151) | 0.076 |
| Observations | 893 | 893 | ||
| Pseudo | 0.015 | 0.005 | ||
| Deviance | 1879.29 | 1931.85 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -939.65 | -965.93 | ||
| Akaike Inf. Crit. | 1891.29 | 1943.85 | ||
| Ordinary Least Squares (OLS) Regression | Dependent variable | |
| Coefficient | ||
| Co-partisan × Neutral | -0.025 (0.017) | 0.133 |
| Cross-partisan × Argue | -0.097 (0.017) | <0.001 |
| Co-partisan × Argue | -0.103 (0.017) | <0.001 |
| Cross-partisan × Support | -0.051 (0.017) | 0.002 |
| Co-partisan × Support | -0.068 (0.017) | <0.001 |
| Constant | 0.873 (0.008) | <0.001 |
| Observations | 444 | |
| 0.118 | ||
| Adjusted | 0.108 | |
| Residual Std. Error | 0.100 (df = 438) | |
| F Statistics | 11.674 (df = 5; 438) | |
| Ordinary Least Squares (OLS) Regression | Dependent variable | |
| Coefficient | ||
| Co-partisan × Neutral | -0.014 (0.016) | 0.381 |
| Cross-partisan × Argue | -0.107 (0.016) | <0.001 |
| Co-partisan × Argue | -0.106 (0.015) | <0.001 |
| Cross-partisan × Support | -0.066 (0.016) | <0.001 |
| Co-partisan × Support | -0.061 (0.016) | <0.001 |
| Constant | 0.891 (0.011) | <0.001 |
| Observations | 449 | |
| 0.157 | ||
| Adjusted | 0.147 | |
| Residual Std. Error | 0.096 (df = 443) | |
| F Statistics | 16.481 (df = 5; 443) | |
| Ordinal Logistical Regression | Dependent variable | Dependent variable | ||
| odds ratio | odds ratio | |||
| 2.643 (0.987) | 0.007 | 1.534 (1.001) | 0.125 | |
| Feedback (Argue) | -1.210 (1.362) | 0.374 | -1.944 (1.346) | 0.148 |
| Feedback (Support) | -0.583 (1.228) | 0.635 | -0.765 (1.238) | 0.537 |
| Feedback (Argue) × | 1.070 (1.684) | 0.525 | 2.282 (1.663) | 0.170 |
| Feedback (Support) × | 1.019 (1.463) | 0.486 | 0.703 (1.471) | 0.633 |
| Observations | 893 | 893 | ||
| Pseudo | 0.027 | 0.011 | ||
| Deviance | 1857.37 | 1921.71 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -928.68 | -960.86 | ||
| Akaike Inf. Crit. | 1871.37 | 1935.71 | ||
| Ordinal Logistical Regression | Dependent variable | Dependent variable | ||
| odds ratio | odds ratio | |||
| 3.233 (0.623) | <0.001 | 2.443 (0.620) | <0.001 | |
| Feedback (Argue) × | -0.410 (0.204) | 0.045 | -0.104 (0.202) | 0.607 |
| Feedback (Support) × | 0.349 (0.187) | 0.062 | -0.167 (0.185) | 0.366 |
| Observations | 893 | 893 | ||
| Pseudo | 0.027 | 0.009 | ||
| Deviance | 1858.17 | 1923.81 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -929.08 | -961.9 | ||
| Akaike Inf. Crit. | 1868.17 | 1933.81 | ||
| Ordinal Logistical Regression | Dependent variable | Dependent variable | ||
| odds ratio | odds ratio | |||
| Feedback (Neutral) × | 3.233 (0.623) | <0.001 | 2.443 (0.620) | <0.001 |
| Feedback (Argue) × | 2.823 (0.701) | <0.001 | 2.339(0.698) | <0.001 |
| Feedback (Support) × | 3.582 (0.670) | <0.001 | 2.276 (0.660) | <0.001 |
| Observations | 893 | 893 | ||
| Pseudo | 0.027 | 0.009 | ||
| Deviance | 1858.17 | 1923.81 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -929.08 | -961.9 | ||
| Akaike Inf. Crit. | 1868.17 | 1933.81 | ||
| Ordinal Logistical Regression | Dependent variable | Dependent variable | ||
| odds ratio | odds ratio | |||
| 3.240 (0.648) | <0.001 | 2.317 (0.644) | <0.001 | |
| Feedback (Argue) × | -0.331 (0.166) | 0.046 | -0.117 (0.164) | 0.477 |
| Feedback (Support) × | 0.272 (0.157) | 0.083 | -0.154 (0.155) | 0.323 |
| Observations | 893 | 893 | ||
| Pseudo | 0.027 | 0.09 | ||
| Deviance | 1858.01 | 1923.61 | ||
| Null Deviance | 1908.84 | 1942.18 | ||
| Log likelihood | -929 | -961.81 | ||
| Akaike Inf. Crit. | 1868.01 | 1933.61 | ||

### v3

@@
 In response, these platforms have implemented various content moderation strategies.
-One such method, Community Notes on X, relies on crowdsourced fact-checking and has gained traction, though it faces challenges such as partisan bias and delays in verification.
-This study explores an AI-assisted hybrid moderation framework in which participants receive AI-generated feedback —supportive, neutral, or argumentative —on their notes and are asked to revise them accordingly.
+One such method, Community Notes (formerly Birdwatch) on X (formerly Twitter), relies on crowdsourced fact-checking and has gained traction.
+However, it faces challenges such as partisan bias and delays in verification.
+This study explores an AI-assisted hybrid moderation framework in which participants receive AI-generated feedback, supportive, neutral, or argumentative, on their notes and are asked to revise them accordingly.
 The results show that incorporating feedback improves the quality of notes, with the most substantial gains resulting from argumentative feedback.
@@
 Introduction
-The emergence of social media platforms has transformed how users consume and share information.
+The emergence of social media has transformed how users consume and share information.
 Studies show that these platforms have become primary news sources for many individuals [1, 2] and are widely used by politicians as campaign tools [3].
 This shift presents challenges for platform owners in monitoring and verifying the information being shared.
-The widespread dissemination of misleading content and misinformation has become a primary concern for social media platforms.
+The widespread dissemination of misleading content and misinformation has become a primary concern for social media.
 This issue is especially pronounced during significant socio-political events, such as the 2016 U.S.
@@
 A 2020 report found that 90% of Republicans in the United States believed social media platforms censor content based on political views [11].
-Community Notes rely on the wisdom of the crowd to identify misleading content and leverage the diversity of users on the platform to produce moderated outcomes.
-The process involves displaying specific posts to users in the Community Notes program and asking them to write a note that provides context for those posts.
-Users are also asked to indicate whether they believe the post is misleading and to provide reasoning for their judgment.
-Subsequently, other users rate the helpfulness of these notes.
-Initially, notes with high helpfulness ratings appeared beneath the posts in users’ timelines.
+In contrast, Community Notes rely on the wisdom of the crowd to identify misleading content and leverage the diversity of users on the platform to produce moderated outcomes.
+Users write a note on a misleading post to provide additional context.
+Other users can then evaluate these notes by rating their helpfulness.
+In the pilot version, notes with high helpfulness ratings appeared beneath the posts in users’ timelines.
 However, early analyses revealed strong partisan behaviour.
@@
 In a recent example, De et al.
-(2025) fine-tuned an LLM on the Community Notes dataset to generate what they call “supernotes,” which were rated significantly more helpful than notes written by Community Notes users for the same posts [30].
+(2025) fine-tuned an LLM on the Community Notes dataset to generate what they call “supernotes”, which were rated significantly more helpful than notes written by Community Notes users for the same posts [30].
+Similarly, recent work has shown that AI can extract fact-checking information from user comments on a post to produce contextual notes that are rated more helpful than Community Notes [31].
 These findings suggest that LLMs are not only capable of understanding complex content but may also outperform human-written moderation in certain cases.
@@
 This hybrid approach leverages the complementary strengths of humans and machines, enabling moderation that is both more nuanced and scalable [35].
+AI systems can process vast volumes of content at speeds unattainable for human moderators.
+Humans, in turn, provide contextual understanding, moral reasoning, and sensitivity to cultural and political nuances that remain challenging for AI to replicate.
 Schmitt et al.
@@
 To mitigate these risks, such systems must be carefully designed to avoid unintentionally amplifying bias and division on social media platforms.
-Despite these risks, political diversity itself can be a valuable asset in improving community-based moderation efforts.
+One proposed strategy is the deliberate introduction of “biased AI”.
+Typically, AI assistants are trained to generate outputs that users rate highly [23, 40].
+However, in hybrid systems where both humans and AI are intended to contribute their respective strengths, overly compliant AI may lead users to over-rely on the system and engage less of their own cognitive resources.
+This phenomenon, referred to as metacognitive laziness, has been observed across different tasks [41, 42].
+Prior research shows that strategic biases can reactivate human critical thinking and improve decision-making [43].
+Introducing bias in controlled ways may enhance human productivity within Human-AI systems.
+For instance, a recent study on collaboration between humans and biased AI in evaluating the veracity of news headlines found that participants performed better when exposed to opposing views.
+However, they regarded the partisan AI as less trustworthy [44].
+Furthermore, introducing AI with perspectives that differ from those of human collaborators can inject political diversity into the moderation process.
+Political diversity itself can be a valuable asset in improving community-based moderation efforts.
 One of the key strengths of Community Notes is its ability to harness ideological diversity—even polarisation—to produce more balanced and neutral notes [45].
@@
 Moreover, studies suggest that exposure to cross-partisan messages can reduce both affective and issue-based polarisation [50, 51].
-These findings indicate that, when properly structured, political diversity not only does not hinder collaboration but also can even enhance it.
-In a recent study, researchers paired two participants to collaboratively write Community Notes.
+These findings suggest that, when properly structured, political diversity not only does not hinder collaboration but can also enhance it.
+In a recent study, researchers paired two participants to write Community Notes collaboratively.
 They found that duos produced more helpful notes than individuals, and that politically diverse groups were especially effective when evaluating posts with Republican-leaning content [52].
 These results further support the idea that political diversity can strengthen the note-writing process in Community Notes, provided it is applied in a way that is both effective and efficient.
+While diversity can strengthen the collective intelligence of teams, its benefits depend on how effectively members can collaborate.
+Collective intelligence emerges not only from the diversity of perspectives but also from the team’s capacity to integrate resources, combine individual knowledge into joint problem-solving, and implement solutions in practice [53, 54].
+Research shows that one of the strongest predictors of collective intelligence is the quality of collaboration processes among team members [55].
+These processes include effective interaction mechanisms such as coordination and cooperation, which are essential for enabling groups to reach outcomes that exceed individual contributions [56].
+Importantly, this principle applies not only to human teams but also to hybrid human–AI collaborations [22].
 Building on these insights, this study proposes a human-AI collaboration framework that uses AI-generated feedback to simulate political diversity in content moderation.
@@
 Our focus is on how users revise their notes in response to different types of AI-generated feedback.
+This work contributes to the growing body of research on the dynamics of Human–AI collaboration, especially in sensitive domains such as content moderation.
 Specifically, this study addresses the following research questions:
@@
 Participants () were shown a post and asked to write a note providing additional context.
-Each note was then submitted to a GPT-4 language model, which was prompted to generate one of three types of feedback: argumentative, supportive, or neutral.
+Each note was then submitted to the GPT-4 language model, which was prompted to generate one of three types of feedback: argumentative, supportive, or neutral.
 Argumentative feedback offers counterpoints to the original note, supportive feedback reinforces the note’s arguments, and neutral feedback simply restates its content.
-While all feedback was generated by the AI model, we randomly told some participants that the feedback was written by a human expert.
+We intentionally did not prompt the LLM to adopt a partisan stance (i.e., Democrat or Republican); instead, the prompt focused solely on the arguments in the note, thereby avoiding stereotyping of either the AI or the participants.
+While the language model generated all the feedback, we randomly instructed about half of the participants in each group that a human expert had written the feedback.
 This manipulation allows us to examine the roles of human favouritism [57] and automation bias [37] in shaping participants’ responses.
-Participants were then asked to revise their notes in light of the feedback.
+Participants were then asked to revise their notes in light of the feedback they received.
 In a second experiment (), we crowdsourced evaluations of note quality before and after revision.
@@
 Figure 1 provides a schematic overview of the experimental workflow.
+The argumentative AI is designed to function as a form of “biased AI”, providing counterarguments to the participants’ initial notes.
+This design aimed to encourage participants to leverage their full cognitive resources and to consider opposing viewpoints.
+Consistent with prior research [44], participants were generally less likely to engage with argumentative AI feedback.
+However, among participants who did engage with the AI feedback, argumentative feedback yielded the highest chance of improvement.
+Overall, our results highlight the potential of AI to assist humans in sensitive tasks such as content moderation.
+They further illuminate the mechanisms of collective hybrid intelligence, demonstrating that the productivity of Human-AI teams depends critically on the extent to which humans actively engage with the AI and the kind of feedback AI provides.
+Additionally, our findings suggest that Human-AI collaboration can effectively simulate political diversity in the note-writing process, thereby enhancing the overall performance of Community Notes.
 Results
@@
 The dependent variables capture whether the revised note has declined, was unchanged, or improved according to the Democrat raters () and Republican raters ().
-The results (Table 1) show that is a strong and consistent predictor of note improvement as perceived by both partisan groups.
+The results (Figure 3) show that is a strong and consistent predictor of note improvement as perceived by both partisan groups.
 A comparison of odds ratios further indicates that has a slightly stronger effect on improvement ratings from Democrats than from Republicans.
-| Ordinal Logistical Regression | Dependent variable: | Dependent variable: | ||
-| odds ratio | odds ratio | |||
-| 3.581(0.604) | <0.001 | 2.498(599) | <0.001 | |
-| Observations | 893 | 893 | ||
-| Pseudo | 0.019 | 0.009 | ||
-| Deviance | 1872.86 | 1924.63 | ||
-| Null Deviance | 1908.84 | 1942.18 | ||
-| Log likelihood | -936.43 | -962.32 | ||
-| Akaike Inf. Crit. | 1878.86 | 1930.63 | ||
+The result of the regression is available in Table S1.
 To better understand the factors that influence this engagement metric, we analyse the relationship between and three key variables: the type of post (supporting either Democratic or Republican ideology), the participant’s self-identified partisanship, and the type of feedback received.
-The results from the Ordinary Least Squares (OLS) regression are shown in Table 2.
+The results from the Ordinary Least Squares (OLS) regression are shown in Figure 4 (a).
 The analysis shows that users who received argumentative feedback were less likely to incorporate it into their final note than those who received supportive feedback.
 Additionally, Democratic participants were more likely than Republicans to use the feedback in their final notes.
-However, the type of post did not have a significant effect on this behaviour.
+However, the type of post did not have a significant effect on this behaviour (see Table S2).
 However, analysing the type of post in isolation may not provide the complete picture.
@@
 To investigate this, we introduce a binary alignment variable that captures the relationship between a participant’s self-identified partisanship and the post’s stance: co-partisan if the participant and the post share the same ideology, and cross-partisan if they do not.
-A post is classified as co-partisan when it supports Democrats (Republicans) and the participant identifies as a Democrat (Republican); otherwise, it is classified as cross-partisan.
-| Ordinary Least Squares (OLS) Regression | Dependent variable: | |
-| Coefficient | ||
-| Feedback (Argue) | -0.094(0.008) | <0.001 |
-| Feedback (Support) | -0.052 (0.008) | <0.001 |
-| Post (D) | 0.001 (0.007) | 0.880 |
-| Participant (D) | 0.021 (0.007) | 0.001 |
-| Constant | 0.850 (0.007) | <0.001 |
-| Observations | 893 | |
-| 0.143 | ||
-| Adjusted | 0.139 | |
-| Residual Std. Error | 0.097 (df = 888) | |
-| F Statistics | 36.899 (df = 4; 888) | |
-To examine how political alignment affects engagement with feedback, we conducted an OLS regression with as the dependent variable, including an interaction term between co-partisanship and feedback type.
-The results of this analysis are shown in Table 3.
-The findings indicate that when participants share the same political affiliation as the post, they are less likely to incorporate argumentative feedback.
-A similar pattern emerges for supportive feedback: participants are also less likely to adopt supportive feedback language when the post aligns with their political views.
-| Ordinary Least Squares (OLS) Regression | Dependent variable | |
-| Coefficient | ||
-| Co-partisan × Neutral | -0.019 (0.011) | 0.090 |
-| Cross-partisan × Argue | -0.102 (0.012) | <0.001 |
-| Co-partisan × Argue | -0.104 (0.011) | <0.001 |
-| Cross-partisan × Support | -0.059 (0.011) | <0.001 |
-| Co-partisan × Support | -0.064 (0.011) | <0.001 |
-| Constant | 0.871 (0.008) | <0.001 |
-| Observations | 893 | |
-| 0.135 | ||
-| Adjusted | 0.130 | |
-| Residual Std. Error | 0.098 (df = 887) | |
-| F Statistics | 27.711 (df = 5; 887) | |
+A post is classified as co-partisan when it supports Democrats (or Republicans) and the participant identifies as a Democrat (or Republican); otherwise, it is classified as cross-partisan.
+To examine how political alignment affects engagement with feedback, we conducted an OLS regression using as the dependent variable, including an interaction term between co-partisanship and feedback type.
+The results, shown in Figure 4 (b), reveal no significant differences in how users incorporate feedback into their notes, regardless of whether they share the same political affiliation as the post.
+This holds for both supportive and argumentative feedback (See Table S3).
 The second treatment condition manipulated the labelled source of the feedback.
@@
 To explore this, we replicate the previous regression analysis on two separate subsets of the data: one in which feedback was attributed to an AI agent, and another in which it was attributed to a human expert.
-Figure 3 presents the marginal predicted values of across the different treatment conditions for both feedback source labels.
+Figure 5 presents the values of across the different treatment conditions for both feedback source labels.
 While the overall patterns are consistent across conditions, we observe a slight increase in when the feedback is labelled as AI-generated, specifically for cross-partisan posts.
@@
 These results suggest a marginal preference for AI-labelled feedback when participants engage with content that opposes their political views.
-However, the differences are not statistically significant (e.g., predicted for AI in cross-partisan × argumentative = 0.777 vs.
-0.762 for human expert, ; predicted for AI in cross-partisan × supportive = 0.822 vs.
+However, the differences are not statistically significant (e.g., for AI in cross-partisan × argumentative = 0.777 vs.
+0.762 for human expert, ; for AI in cross-partisan × supportive = 0.822 vs.
 0.803 for human expert, ), indicating that the labelled source of feedback does not meaningfully influence user behaviour in this context.
-To construct the full model and evaluate the effect of feedback type conditional on participant engagement, we estimated an ordinal logistic regression including both the main effects of feedback type and its interaction with .
+To construct the complete model and evaluate the effect of feedback type, conditional on participant engagement, we estimated an ordinal logistic regression that included both the main effects of feedback type and its interaction with .
 The dependent variables are and , which reflect the improvement of the notes according to Democrat and Republican raters, respectively.
-This model allows us to assess whether argumentative or supportive feedback is more likely to lead to improvements in perceived helpfulness when participants actively engage with the feedback, and whether these effects differ across political subgroups.
-The results of this analysis are presented in Table 4.
+This model enables us to determine whether argumentative or supportive feedback is more likely to lead to improvements in perceived helpfulness when participants actively engage with the feedback, and whether these effects vary across different political subgroups.
+The results of this analysis are presented in Figure 6.
 The findings indicate that once participants incorporate feedback into their notes, both argumentative and supportive feedback increase the likelihood of producing a more helpful note, as rated by both Democratic and Republican evaluators.
 However, argumentative feedback shows a larger effect size than supportive feedback, suggesting that when participants engage with feedback that challenges their views and includes counterarguments, the resulting notes are more likely to be of higher quality.
-| Ordinal Logistical Regression | Dependent variable | Dependent variable | ||
-| odds ratio | odds ratio | |||
-| Feedback (Argue) | -1.210 (1.362) | 0.374 | -1.944 (1.346) | 0.148 |
-| Feedback (Support) | -0.583 (1.228) | 0.635 | -0.765 (1.238) | 0.536 |
-| Feedback (Neutral) × | 2.643 (0.987) | 0.007 | 1.534 (1.001) | 0.125 |
-| Feedback (Argue) × | 3.713 (1.370) | 0.006 | 3.817 (1.331) | 0.004 |
-| Feedback (Support) × | 3.662 (1.086) | <0.001 | 2.237 (1.079) | 0.038 |
-| Observations | 893 | 893 | ||
-| Pseudo | 0.027 | 0.011 | ||
-| Deviance | 1857.37 | 1921.71 | ||
-| Null Deviance | 1908.84 | 1942.18 | ||
-| Log likelihood | -928.68 | -960.86 | ||
-| Akaike Inf. Crit. | 1871.37 | 1935.71 | ||
 Discussion
-As Community Notes gains traction across major social media platforms as a leading approach to crowd-based content moderation [8], concerns have been raised about its efficiency and long-term sustainability [13, 48, 15].
-One emerging response to these challenges is the integration of generative AI—particularly large language models (LLMs)—into the Community Notes ecosystem.
+As Community Notes gains traction across major social media platforms as a leading approach to crowd-based content moderation [8, 58], concerns have been raised about its efficiency and long-term sustainability [13, 59, 15, 60].
+One emerging response to these challenges is the integration of generative AI, particularly large language models (LLMs), into the Community Notes ecosystem.
 In this study, we introduce a framework that leverages GPT-4 to provide users with different types of feedback on their notes.
-This framework is designed both to evaluate the effectiveness of AI-assisted feedback and to deepen our understanding of the dynamics of human–AI collaboration in this context.
+This framework is designed both to evaluate the effectiveness of AI feedback and to deepen our understanding of the dynamics of human–AI collaboration in this context.
 We found that within this framework, the effectiveness of feedback in improving note quality depends not only on its type but also on the extent to which participants engage with it.
-Studies have shown that two key factors in collective intelligence are the effort a group invests and how effectively members leverage each other’s knowledge and skills [49].
-Our findings align with this: participants who engaged with AI-generated feedback more and incorporated it into their revisions produced better perceived notes.
-Specifically, the Feedback Acceptance rate ()—which reflects how closely participants’ final notes match the feedback—emerged as a significant predictor of note improvement for both Democratic and Republican raters ( and ).
-A high indicates strong engagement with the feedback, while a low often corresponds to users disregarding the feedback entirely, sometimes resulting in notes of lower quality than the original (See Figure SI10).
-This behaviour mirrors patterns observed on social media, where users often react defensively to content that challenges their political beliefs [50].
-Our findings suggest that when participants move beyond this reflexive partisan response and actively engage with feedback, they produce higher-quality notes.
-Among participants who actively engaged with the feedback, our findings show that argumentative feedback was more likely to lead to improvements in note quality.
+Studies have shown that two essential factors in collective intelligence are the effort a group invests and how effectively members leverage each other’s knowledge and skills [61].
+Our findings suggest that human-AI collective efforts follow the same principle: participants who engaged with AI-generated feedback more and incorporated it into their revisions produced notes that were perceived as better.
+Specifically, the Feedback Acceptance rate (), which reflects how closely participants’ final notes match the feedback, emerged as a significant predictor of note improvement for both Democratic and Republican raters ( and ).
+A high indicates strong engagement with the feedback, while a low often corresponds to users disregarding the feedback entirely, sometimes resulting in notes of lower quality than the original (See Figure S10).
+is lower once the feedback is argumentative; This pattern mirrors behaviours observed on social media, where users often respond defensively to content that challenges their political beliefs [62], and aligns with prior research showing that users distrust AI systems offering opposing viewpoints [44].
+Nevertheless, our findings suggest that when participants move beyond this reflexive partisan response and actively engage with feedback, they produce higher-quality notes.
+Among participants who actively engaged with the feedback, our findings indicate that argumentative feedback was more likely to result in improvements in note quality.
 This may be because engaging with counterarguments encourages users to reconsider and refine their positions, resulting in notes that are more balanced, better justified, and ultimately more helpful.
 Prior research suggests that in problem-solving contexts, when AI systems are designed to complement human cognitive biases, they can prompt users to discover novel solutions, thereby improving outcomes [63].
+Similarly, a biased AI assistant can motivate humans to challenge competing views and increase their cognitive effort [44].
+In our study, the argumentative AI functioned as such a biased assistant, providing an opposing perspective that nudged participants to consider alternative viewpoints and thereby introduced diversity into the note-writing process.
 Relatedly, Juncosa et al.
@@
 Our current findings suggest that argumentative AI feedback may function as a form of political diversity, exposing users to alternative viewpoints and, in doing so, fostering the production of higher-quality content.
-In our second research question, we examine how the political alignment between participants and posts affects the perceived helpfulness of the notes.
-The results indicate that when participants and posts share the same political ideology (i.e., are co-partisan), participants are less likely to incorporate feedback into their notes, reducing the likelihood of producing a higher-quality note.
+In our second research question, we investigated how the political alignment between participants and posts affects the perceived helpfulness of notes.
+Our findings indicate no difference in how users incorporate feedback when the post aligns with their political ideology (co-partisan) versus when it does not (cross-partisan).
 This pattern holds for both argumentative and supportive feedback.
-One possible explanation for this behaviour is that when participants encounter a post aligned with their own beliefs, they may become less open to revision, regardless of the content or tone of the feedback [52, 53].
+Because the Feedback Acceptance rate () is unaffected by participant-post alignment, note improvement is similarly unaffected.
 Lastly, in our third research question, we examined how the labelled source of feedback, whether an AI agent or a human expert, influences participant behaviour and, consequently, the quality of their notes.
-We find a slight preference for AI-labelled feedback when the post comes from the opposing political party (i.e., in cross-partisan contexts).
+We find a slight preference for AI-labelled feedback when the post comes from the opposing political party (i.e., in cross-partisan contexts), though this difference was not statistically significant.
 In these cases, participants are more likely to engage with feedback they believe was generated by an AI agent.
@@
 We demonstrate that integrating a large language model (LLM) into the Community Notes system can enhance its efficiency, provided the design preserves human agency and the platform’s core participatory nature.
-In our framework, the AI functions solely as a reviewer offering suggestions, while users retain full control over the content of their notes, helping to prevent the imposition of any algorithmic bias.
+In our framework, the AI functions solely as a reviewer offering suggestions.
+At the same time, users retain complete control over the content of their notes, helping to prevent the imposition of any algorithmic bias.
 A key contribution of our approach is the introduction of political diversity at the note-writing stage, rather than relying solely on diverse raters at the evaluation stage.
@@
 This suggests that such feedback not only promotes quicker production of helpful notes but also encourages users to consider counterarguments and revise their contributions in a more balanced and thoughtful way, ultimately discouraging the reinforcement of partisan narratives.
-While our findings highlight the potential of integrating argumentative AI feedback into Community Notes, there are important limitations to consider.
-First, we do not assess the objective truthfulness of the notes.
-Instead, note quality in this study is evaluated through crowdsourced ratings, with groups of Democrats and Republicans judging notes based on their perceived helpfulness.
-This method does not provide a direct measure of factual accuracy—for instance, a professional fact-checker might reach different conclusions.
-However, our approach reflects the current practice in Community Notes, where community members determine which notes are approved and displayed beneath posts.
-While some researchers question the reliability of crowd-based methods for identifying misinformation, it is important to note that the goal of Community Notes is not necessarily to debunk falsehoods, but to add context to potentially misleading posts.
-These posts may not contain explicit inaccuracies, but can still lead readers to incorrect conclusions due to missing or unbalanced information.
-In such cases, a politically diverse and balanced group of laypeople may be well positioned to assess whether the added context is informative and helpful.
-Second, findings from an experimental setting may not fully generalise to real-world contexts.
-For example, participants in Community Notes are volunteers who may be motivated by political or ideological beliefs, whereas participants in our study were recruited and compensated for their involvement.
+Compared to pairing two politically diverse human participants to co-author a note, this approach is also less resource-intensive.
+However, realising the benefits of such hybrid collective intelligence requires the human participants to actively engage with the AI, much like the way collective intelligence emerges in human groups.
+Our findings shed light on the underlying mechanisms through which hybrid collective intelligence emerges in human–AI collaboration.
+On the one hand, engagement with AI feedback is crucial.
+Without active engagement, the benefits of AI feedback cannot be realised, just as collective intelligence among humans only works when they collaborate.
+On the other hand, the argumentative nature of the feedback plays a central role in fostering political diversity and balance at the note-writing stage.
+By prompting participants to confront counterarguments and revise their notes accordingly, argumentative feedback encourages deeper deliberation and reduces the risk of partisan echo chambers.
+Together, these mechanisms illustrate how hybrid collectives can leverage AI not as a replacement for human reasoning, but as a catalyst that enhances the quality and inclusiveness of collective outputs.
+While our findings highlight the potential of integrating argumentative AI feedback into Community Notes, several important limitations must be acknowledged.
+First, this study does not assess the objective factual accuracy of the notes.
+Instead, note quality is evaluated through crowdsourced ratings, with groups of Democrats and Republicans assessing perceived helpfulness.
+This approach does not provide a direct measure of truthfulness, for example, professional fact-checkers might arrive at different conclusions.
+However, our method aligns with the existing structure of Community Notes, where note approval and display are determined by community consensus rather than expert adjudication.
+Augenstein et al.
+(2025) [60] describe this as an epistemological distinction between traditional fact-checking and Community Notes, where traditional fact-checking seeks objective truth verified by experts, whereas Community Notes reflect subjective judgments based on users’ perspectives.
+Importantly, the aim of Community Notes is not solely to debunk falsehoods but to add context to potentially misleading posts.
+These posts may not contain outright inaccuracies, yet still risk misinforming readers through omission or imbalance.
+In such cases, a politically diverse and balanced group of laypeople is well-suited to assess whether the contextual information provided is genuinely informative and helpful [64].
+Studies have found that there is substantial agreement with the Community Notes and professional fact checkers [65].
+Second, findings derived from an experimental setting may not fully generalise to real-world contexts.
+For instance, Community Notes participants are volunteers often motivated by political or ideological commitments, whereas participants in our study were recruited and compensated.
+This difference is reflected in the tone and structure of the notes produced here, which diverge from those typically found on the Community Notes platform.
+Additionally, in community-based efforts, a small group of ”power users” often comes to dominate content creation, developing a specialised language and tone tailored to the platform’s norms.
+This phenomenon is evident in Community Notes, where the widespread use of shorthand, such as “NNN” (“No Note Needed”), signals that a post is not misleading and does not require annotation.
+New or casual users are unlikely to be familiar with such conventions, resulting in experienced contributors acting as gatekeepers, a dynamic that may hinder platform scalability [66].
+It is essential to emphasize that this study did not aim to replicate Community Notes, but rather to investigate the impact of integrating AI-driven argumentative assistance into the community-based content moderation process.
 Third, the “breaking news problem” remains a persistent challenge.
+This occurs when a note addresses an event of which the AI is unaware.
 For AI to provide meaningful feedback, it must have a sufficient understanding of the topic.
-When this is lacking, the AI may still offer feedback that encourages critical thinking, effectively acting as a nudging algorithm [54].
+When this is missing, the AI may still add value by prompting users to think critically—acting like a nudging algorithm [67] that encourages them to seek out reliable sources and investigate further.
 That said, recent advancements in large language models have enabled web-search capabilities, which may allow them to engage more effectively with current events in future applications.
@@
 However, the presence of AI alone is insufficient to ensure effective collaboration.
-Its impact depends on active user engagement, which is shaped by factors such as the participant’s partisanship and the political alignment of the post.
+Its impact depends on active user engagement.
 Future work could extend these insights by examining longer-term interactions between users and AI to understand how engagement patterns evolve over time.
@@
 Data used in the study are available on Harvard Dataverse (doi:10.7910/DVN/SWTACT).
+Declaration of generative AI in scientific writing
+During the preparation of this work, the authors used ChatGPT 4.0 to improve the writing style of this article.
+After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
 Acknowledgments
@@
 After providing consent and entering their Prolific ID, participants were asked to indicate their political partisanship on a nine-point scale ranging from strongly Democrat to strongly Republican.
-A screenshot of this page is shown in Figure SI1.
+A screenshot of this page is shown in Figure S1.
 Participants who selected the neutral midpoint option were excluded from the experiment.
@@
 At this stage, they were informed that the feedback they would receive could come either from a human expert or an artificial intelligence machine.
-Figure SI2 shows a screenshot of this page.
+Figure S2 shows a screenshot of this page.
 In the first task of the experiment, participants were instructed to read a social media post and write a note aimed at providing additional context and assessing whether the post was misleading or not.
@@
 Participants were given 10 minutes to complete this task.
-Figure SI3 shows an example of the task page.
-After completing the first task, participants were shown a “please wait” page, which informed them about the source of the upcoming feedback.
+Figure S3 shows an example of the task page.
+After completing the first task, participants were shown a ”please wait” page, which informed them about the source of the upcoming feedback.
 They remained on this page for 2.5 minutes to create the impression that the feedback might be provided by a human.
-Figure SI4 displays two screenshots of this page, illustrating both feedback source conditions.
+Figure S4 displays two screenshots of this page, illustrating both feedback source conditions.
 Afterwards, participants were given seven minutes to review the feedback and revise their notes accordingly.
@@
 Additionally, participants were reminded of the feedback source.
-An example of this page is shown in Figure SI5.
+An example of this page is shown in Figure S5.
 After the second task, participants completed a post-experiment survey where they rated their comfort level with various technologies.
-Figure SI6 shows the survey page displaying the full list of questions.
+Figure S6 shows the survey page displaying the full list of questions.
 Participants were given four minutes to complete the survey.
 Subsequently, participants were asked to identify the source of the feedback as an attention check.
-Those who selected an incorrect option or chose “I do not remember” were excluded from the dataset.
+Those who selected an incorrect option or chose ”I do not remember” were excluded from the dataset.
 Finally, on the debrief page, participants were informed that all feedback was generated by an AI agent and that no human experts were involved.
@@
 Afterwards, participants were asked to indicate their partisanship on a nine-point scale, consistent with the note-writing experiment.
-Those who selected the “neutral” option were excluded from the study.
+Those who selected the ”neutral” option were excluded from the study.
 In the main task of the experiment, participants were shown a social media post accompanied by 10 notes written about that post.
@@
 Participants had 7 minutes to evaluate all 10 notes.
-Figure SI7 shows a screenshot of the task description.
+Figure S7 shows a screenshot of the task description.
 Each participant evaluated two different social media posts, each with its own set of 10 notes.
@@
 Neutral: Read the following tweet:tweet. Read the comment made in response to the tweet: comment. Rephrase the comment. Do not add any evidence or contradict the comment in any way.
-The ”tweet” refers to the content of the post, while the ”comment” is created by appending the sentence “The tweet is misleading” or “The tweet is not misleading” to the note written by the participant in the first task.
+The ”tweet” refers to the content of the post, while the ”comment” is created by appending the sentence ”The tweet is misleading” or ”The tweet is not misleading” to the note written by the participant in the first task.
 This sentence is selected based on the participant’s classification of the post.
@@
 Additional Tables
+| Ordinal Logistical Regression | Dependent variable: | Dependent variable: | ||
+| odds ratio | odds ratio | |||
+| 3.581(0.604) | <0.001 | 2.498(0.599) | <0.001 | |
+| Observations | 893 | 893 | ||
+| Pseudo | 0.019 | 0.009 | ||
+| Deviance | 1872.86 | 1924.63 | ||
+| Null Deviance | 1908.84 | 1942.18 | ||
+| Log likelihood | -936.43 | -962.32 | ||
+| Akaike Inf. Crit. | 1878.86 | 1930.63 | ||
+| Ordinary Least Squares (OLS) Regression | Dependent variable: | |
+| Coefficient | ||
+| Feedback (Argue) | -0.094(0.008) | <0.001 |
+| Feedback (Support) | -0.052 (0.008) | <0.001 |
+| Post (D) | 0.001 (0.007) | 0.880 |
+| Participant (D) | 0.021 (0.007) | 0.001 |
+| Constant | 0.850 (0.007) | <0.001 |
+| Observations | 893 | |
+| 0.143 | ||
+| Adjusted | 0.139 | |
+| Residual Std. Error | 0.097 (df = 888) | |
+| F Statistics | 36.899 (df = 4; 888) | |
+| Ordinary Least Squares (OLS) Regression | Dependent variable | |
+| Coefficient | ||
+| Co-partisan × Neutral | -0.019 (0.011) | 0.090 |
+| Cross-partisan × Argue | -0.102 (0.012) | <0.001 |
+| Co-partisan × Argue | -0.104 (0.011) | <0.001 |
+| Cross-partisan × Support | -0.059 (0.011) | <0.001 |
+| Co-partisan × Support | -0.064 (0.011) | <0.001 |
+| Constant | 0.871 (0.008) | <0.001 |
+| Observations | 893 | |
+| 0.135 | ||
+| Adjusted | 0.130 | |
+| Residual Std. Error | 0.098 (df = 887) | |
+| F Statistics | 27.711 (df = 5; 887) | |
+| Ordinal Logistical Regression | Dependent variable | Dependent variable | ||
+| odds ratio | odds ratio | |||
+| Feedback (Argue) | -1.210 (1.362) | 0.374 | -1.944 (1.346) | 0.148 |
+| Feedback (Support) | -0.583 (1.228) | 0.635 | -0.765 (1.238) | 0.536 |
+| Feedback (Neutral) × | 2.643 (0.987) | 0.007 | 1.534 (1.001) | 0.125 |
+| Feedback (Argue) × | 3.713 (1.370) | 0.006 | 3.817 (1.331) | 0.004 |
+| Feedback (Support) × | 3.662 (1.086) | <0.001 | 2.237 (1.079) | 0.038 |
+| Observations | 893 | 893 | ||
+| Pseudo | 0.027 | 0.011 | ||
+| Deviance | 1857.37 | 1921.71 | ||
+| Null Deviance | 1908.84 | 1942.18 | ||
+| Log likelihood | -928.68 | -960.86 | ||
+| Akaike Inf. Crit. | 1871.37 | 1935.71 | ||
 | Ordinal Logistical Regression | Dependent variable | Dependent variable | ||